A function is defined as f(x)=x(squared)+3
Find a simplified expression for f(a+2)-f(a-5)
1
Expert's answer
2012-10-23T10:14:14-0400
we can write f(a+2)-f(a-5)=(a+2)^2 +3 - ( (a-5)^2+3)=(a+2)^2 -
(a-5)^2=(a+2+a-5)(a+2 -(a-5) )=(2a-3)*7=14a-21